OpenStudy (cresciez):

Help please? Simple geometry :> A cone has a volume of 5 cubic inches. What is the volume of a cylinder that the cone fits exactly inside of?

OpenStudy (cresciez):

15 in3 20 in3 25 in3 30 in3

OpenStudy (cresciez):

I'm only confused about the formulas. I've researched this question and have the answer, but the formulas for a cone are different. The one I've been given on my homework is V=1/3(pi r^2)(h), but the only way to solve it is by using V=1/3bh. Can someone explain?

OpenStudy (cresciez):

@ganeshie8 I know you might be busy, but I would appreciate it so much if you could clear this up! :>

OpenStudy (cresciez):

@Directrix

OpenStudy (cresciez):

@pooja195

OpenStudy (cresciez):

@jigglypuff314 @Preetha please ;-;

OpenStudy (cresciez):

@freckles @IrishBoy123 please?

OpenStudy (cresciez):

@SamsungFanBoy @ILovePuppiesLol

OpenStudy (cresciez):

Please guys, I have so many more assignments and nobody is bothering to help.

OpenStudy (cresciez):

@LifeIsADangerousGame

OpenStudy (tgannon998):

I can help you in a minute

OpenStudy (cresciez):

Yay! Thanks !

OpenStudy (tgannon998):

np

OpenStudy (tgannon998):

So the cone is one third the size of the cylinder so multiply the 5 times 3 and you get your answer.

OpenStudy (cresciez):

Yes, but as I said before: I'm only confused about the formulas. I've researched this question and have the answer, but the formulas for a cone are different. The one I've been given on my homework is V=1/3(pi r^2)(h), but the only way to solve it is by using V=1/3bh. Can someone explain?

OpenStudy (cresciez):

Is it circumference and volume?

OpenStudy (cresciez):

I think there are two formulas for each shape but I'm not sure...

OpenStudy (tgannon998):

Yes

OpenStudy (cresciez):

Hmm.. well thanks for your help!

OpenStudy (tgannon998):

np

OpenStudy (michaelbp):

do you understand everything